If the larger of two numbers is divided by the smaller, the quotient is 7 and the remainder is 19. But if 3 times the greater is divided by twice the smaller, the quotient is 11 and the remainder is 19. What are the numbers?
Answer by ikleyn(52867)

.
I just solved it for you at this link
https://www.algebra.com/algebra/homework/Finance/Finance.faq.question.1097559.html

https://www.algebra.com/algebra/homework/Finance/Finance.faq.question.1097559.html


For your convenience, I copy that solution here one more time.
I added the check, which shows that the solution is correct.
If you have another answer, it is wrong.

Let x be the larger of the two numbers and y be the smaller.

Then the condition gives you these two equations


x = 7y + 19,           (1)

3x = 11*(2y) + 19.     (2)


Substitute (1) into (2). You will get


3*(7y + 19) = 22y + 19,    or

21y + 57 = 22y + 19  ====>  y = 57 - 19 = 38.


Then  x = 7*y + 19 = 7*38 + 19 = 285.


Answer.  The larger number is 285.  The smaller number is 38.


Check.   Check eq(1):  7*y+19 = 7*38+19 = 285 = x.             ! Correct !

         Check eq(2):  3x = 3*285 = 11*2*38+19 = 11*(2y)+19.   ! Correct !